OpenStudy (anonymous):

I need a bit of help... please determine the quadratic function if the vertex is (1,-4) and the y intercept is -3 answer must be in the form f(x)=ax^2+bx+c

jimthompson5910 (jim_thompson5910):

If the vertex is (1,-4), then (h,k) = (1,-4) meaning that h = 1 and k = -4 y = a(x-h)^2 + k y = a(x-1)^2 + (-4) y = a(x-1)^2 - 4 We're given that it has a y-intercept of -3, so x = 0 and y = -3. Plug these additional values in y = a(x-1)^2 - 4 -3 = a(0-1)^2 - 4 -3 = a(-1)^2 - 4 -3 = a(1) - 4 -3 = a - 4 -3+4 = a 1 = a a = 1 So the function in vertex form is f(x) = 1(x-1)^2 - 4 which can be written as f(x) = (x-1)^2 - 4 Now convert to standard form f(x) = (x-1)^2 - 4 f(x) = x^2-2x+1 - 4 f(x) = x^2-2x-3 So the answer is f(x) = x^2-2x-3

OpenStudy (anonymous):

ahh i see ;)

jimthompson5910 (jim_thompson5910):

that's great :)